OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 9x + 3 and g(x) =(x-3)/9 please help.

OpenStudy (anonymous):

this is what i have so far f(g(x))= 9((x-3)/9)+3 (9x-27)/9 + 3 (9x-24)/9 g(f(x))= ((9x+3)-3)/9 9x/9 does this mean theyre not inverses?

OpenStudy (anonymous):

There's a small mistake you made: \[\begin{align*}f(g(x))&=9\left(\frac{x-3}{9}\right)+3\\ &=\frac{9x-27}{9}+3\\ &=\frac{9x-27}{9}+\color{red}{\frac{27}{9}}\\ &=\frac{9x}{9}\\ &=x \end{align*}\]

OpenStudy (anonymous):

so they are inverses

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you :)